# RadarPLM: Adapting Pre-trained Language Models for Marine Radar Target Detection by Selective Fine-tuning

### v4

@@ Abstract
Recent advances in pre-trained language models (PLMs) have demonstrated their capabilities in capturing universal knowledge, making them promising for radar signal processing applications. Nevertheless, directly fine-tuning PLMs on radar signals is both computationally expensive and prone to overfitting, particularly in low signal-to-clutter ratio (SCR) environments. In this paper, we propose a novel fine-tuning framework for PLM-based marine radar target detection. First, we design a lightweight adaptation module, enabling computationally efficient fine-tuning while preserving the pre-trained model’s general knowledge. Second, a novel preference-aware loss is developed to selectively optimize different feature patches based on their online-evaluated learning values, guiding the model to concentrate on those generalizable feature patterns during optimization. Finally, a binary classification head is retrained based on autoencoder network to further enhance detection performance. Experiments on real-world radar data show that the proposed RadarPLM framework yields at least a 6.35% improvement in detection performance over the existing networks under low SCR conditions. Especially, in small training samples cases, the proposed RadarPLM also achieves significant advantage over existing networks owing to the incorporation of the PLM.
@@ I Introduction
Recently, the research community has experienced a paradigm shift from lightweight, task-specific neural networks toward large-scale foundation models. In particular, pre-trained large language models (LLMs), such as ChatGPT and DeepSeek, have demonstrated strong generalization capabilities across a wide range of complex tasks, including long-form text generation, mathematical reasoning, and code synthesis [12]. Motivated by these advances, an increasing body of research has extended LLMs beyond the textual domain to handle non-textual data modalities, giving rise to multimodal large language models (MLLMs). In parallel, several recent studies [13, 14, 15] have explored the adaptation of pre-trained large language models for time-series analysis and wireless signal processing. These works consistently report notable performance improvements over conventional lightweight task-specific models, particularly in scenarios characterized by limited labeled data. Building upon these developments, this work investigates the feasibility and effectiveness of leveraging PLMs as a powerful backbone network for marine radar signal processing, aiming to enhance model robustness and generalization under challenging data conditions.
However, some challenges exist in adapting PLM for marine radar signal processing applications. First, owing to the large parameter scale of PLMs, their adaptation to marine radar signal processing necessitates lightweight and efficient fine-tuning strategies, particularly for deployment in resource-constrained environments. Such strategies are expected to significantly reduce the number of trainable parameters by several orders of magnitude, while preserving the pre-trained knowledge of PLMs. Second, adapting PLMs to process radar signal features in low SCR conditions often leads to severe overfitting issues. This phenomenon is fundamentally attributed to the intrinsic characteristics of radar signal features:
  • Significant Clutter-Induced Noise: In low SCR environments, small targets are frequently obscured by ocean waves, which inevitably generate mixed-statue segments. The mixed statue segments can cause a low SCR and increase the likelihood of the model overfitting to irrelevant and noisy information.
  • Feature Pattern Convergence Imbalance: Various patterns in different sequence features often exhibit different convergence rates during fine-tuning. For example, simple and regular patterns may rapidly reach convergence during model fine-tuning and then exhibit a tendency for overfitting, whereas other valuable patterns may require more iterations to achieve convergence.
Consequently, uniform optimization across all feature patterns may cause the model to overfit not only to noisy patterns but also to overly simplistic ones, which fail to capture the more generalizable patterns inherent in radar sequence features.
@@
To overcome the above-mentioned issues, including the demand for lightweight fine-tuning and the tendency of overfitting under low SCR conditions, we develop a novel framework named RadarPLM (based on GPT2, an advanced open-source PLM.). The proposed framework effectively bridges the gap between PLMs and radar target detection tasks with a lightweight adaptation module and a selective training strategy. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Subsequently, a radar target detector is built upon GPT2, a highly advanced open-source PLM, and is fine-tuned through a lightweight adaptation module to efficiently align with the specific radar detection task. For model optimization, we design a preference-aware loss function, which is inspired by recent advances [16, 17]. This loss introduces a selective training strategy dynamically adjusts the learning weights of different feature patches according to their online-evaluated learning values, significantly mitigating model overfitting. Finally, the binary classification head is retrained to further improve detection performance.
@@
    We propose a novel framework RadarPLM based on GPT2 model for marine radar signal processing applications. To the best of our knowledge, this work is the first to demonstrate that the recent developed PLMs can serve as a strong starting point for training and optimizing marine radar signal features.
@@
    For fine-tuning, we develop a lightweight adaptation module and design a novel selective training strategy, to reduce finetuning computational overhead and significantly mitigate the risk of model overfitting.
@@
    Extensive experiments validate that RadarPLM achieves new state-of-the-art detection performance on popular marine radar datasets, especially under practical limitations including limited training data and low SCR conditions.
Our preliminary study [18] demonstrated that fine-tuning a pre-trained GPT2 model effectively extracts discriminative information from sequence features, achieving state-of-the-art detection performance. In this paper, we further introduce a lightweight adaptation module and selective training strategy for more lightweight and generalizable adaptation.
@@ II-A Marine Radar Target Detection Based on Deep Learning
II-B Signal Processing Models Based on Pre-trained Language Model
Recently, PLMs such as GPT2 and LLaMA2 have demonstrated exceptional capabilities in signal processing and communication tasks. For example, Zhou et al. [13] propose a unified time series analysis framework by fine-tuning frozen GPT2 [23] and achieving SOTA performance on various datasets. Liu et al. [24] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [25] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [26] propose a multi-task PLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM and Penetrative AI [27, 28] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks.
However, few of these models are specifically devised for radar signal processing applications. Moreover, recent studies [29] have raised significant concerns regarding the universal applicability and direct transferability of PLMs to various signal processing tasks. PLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. In marine radar target detection under low SCR conditions, directly fine-tuning a PLM often leads to severe overfitting issues, limiting its parameter transfer capability. To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
@@ III Proposed Method
(a) IP feature
(b) DSE feature
(c) SMS feature
(d) Amp feature
(e) DP feature (Target)
(f) DP feature (Clutter)
Figure 3: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX database.
@@ III-D1 Methodology
  • Low Importance (Noise/Anomaly): If $\alpha\mathcal{L}_{\theta_{{r}}}$ is large (indicating with high inherent noise), it effectively dampens $s_{b,k}$, even if $\mathcal{L}_{\theta_{{t}}}$ is large, preventing the model from overfitting to noisy or anomalous data.
@@ III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
However, directly optimizing Eq.(16) is computationally prohibitive, as it requires retraining the model for each candidate patch. To enable a practical surrogate, we reformulate this objective using probabilistic decomposition.
B. Derivation of the Surrogate Objective: Applying Bayes’ theorem, the term in Eq.(16) can be decomposed as:
@@
Substituting Eq.(18) into Eq.(17) and taking the logarithm yields:
@@
In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the holdout samples in Eq.(16) can be approximated by the following objective:
@@
D. Conclusion: Consequently, the calculated gap between the target loss $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})$ and the approximated irreducible loss $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})$ (derived from the lightweight reference model) serves as the practical measure of each feature patch’s learning value, which forms the basis for the token weighting strategy defined in Eq.(14).
@@ III-E Autoencoder-Based Binary Classification Head Retraining
As shown in Fig. 1 and Section III-A, the PLM-based network produces token-level outputs of shape $[B,K,2]$, which is inconvenient for final decision. To obtain a binary output for each sample, we therefore retrain an autoencoder-based binary classification head on top of feature representation $F^{\mathrm{LN}}$. It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative representation; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector. The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
@@
    Extract IP, DSE, SMS, Amp, and DP features from the observation vector, respectively, see Eq.(3), (4), (5),(7), and (9).
@@
    Compute the loss for each feature token of the training samples based on the predicted output probabilities of $\theta_{r}$, see Eq.(12).
@@
    Compute the loss for each feature token of the training samples based on the predicted output probabilities of current model, see Eq.(13).
@@
    Fine-tune PLM for target detection via lightweight adaptation module and preference-aware loss function, see Eq.(14) and Eq.(15).
@@
    Train autoencoder-based binary classification head based on the weighted sum of reconstruction loss and CE loss, see Eq.(21), Eq.(22), and Eq.(23).
@@
    Adjust detection threshold $\eta$ for controllable FAR, see Eq.(24).
@@ IV Experiments
Label Name File name Primary Secondary SCR
1 17 19931107_135603_starea 9 8,10,11 Low
2 25 19931108_213827_starea 7 6,8 Low
3 26 19931108_220902_starea 7 6,8 Low
4 54 19931111_163625_starea 8 7,9,10 High
5 280 19931118_023604_stareC0000 8 7,9,10 High
6 283 19931118_035737_stareC0000 10 8,9,11,12 High
7 311 19931118_162658_stareC0000 7 6,8,9 High
8 320 19931118_174259_stareC0000 7 6,8,9 High
@@ IV-A1 Dataset
To ensure sufficient training data, we employ overlapped segmentation following the partition rule in Eq. (1), with parameters set to $M=32$ for target cells and $M=128$ for clutter cells. This process generates 4,079 target samples and more than 9,000 clutter samples per dataset. The samples are divided into three groups: (1) a training set using the first 20% observation time for both the target and clutter cells, (2) a validation set covering 20% to 35% of the observation time, and (3) a test set containing the remaining samples.
@@ IV-A4 Implementation Details
The experimentally relevant hyperparameters and details are shown in Table II. For patching, the size of feature patches is set to 48. For the PLM, the smallest version of GPT2 with $F=768$ feature dimension and the first $L=4$ layers of the GPT2 encoder are deployed. The value of $\alpha$ in Eq.(14) is set to 0.9. All signal feature extraction operations are performed with MATLAB R2016a software. All network training experiments are conducted on a system equipped with an E5-2695v3 CPU, an NVIDIA 3090Ti GPU, and 64 GB of RAM.
IV-B Performance Comparison
(1). HH
(2). HV
(3). VH
(4). VV
Method All Data Low SCR Data
$H$ HV VH $V$ AVG $H$ HV VH $V$ AVG
RNN [10] 46.04 46.66 38.18 17.04 36.98 9.45 13.00 11.67 4.47 9.65
GRU [35] 76.46 77.55 73.96 67.48 73.86 48.99 50.46 49.79 27.33 44.14
OFA [13] 74.19 80.38 79.71 70.36 76.16 49.28 54.18 54.62 26.96 46.26
ADN18 [37] 79.74 83.50 82.01 78.02 80.82 54.76 60.04 31.32 46.11 47.31
RadarPLM 82.51 84.81 84.61 75.90 81.96 60.03 63.19 62.27 44.82 57.58
@@ IV-B Performance Comparison
More Crucially, RadarPLM demonstrates exceptional robustness in the challenging low SCR conditions. As shown in the right section of Table III, while most baselines suffer from severe performance degradation due to strong clutter-induced noise, RadarPLM maintains a remarkable lead. It achieves an average detection rate of 57.58% on low SCR datasets, surpassing the second-best method (ResNet18) by 6.35% and the complex ADN18 method by a significant margin of 10.27%. This result highlights the critical advantage of our method in learning discriminative representation from noisy sequence features.
@@
IV-C Scarce Training Sample Condition
@@
Figure 7: Performance comparison under scarce and abundant training sample conditions. The results show that the RadarPLM approach experiences substantially smaller performance degradation under scarce training sample condition while maintaining the best overall detection performance.
@@
(1). Data1
(2). Data2
(3). Data3
Figure 8: We compare the detection rates achieved by the proposed preference-aware loss against the conventional cross-entropy (CE) loss across various datasets and training epochs. The results demonstrate that while the CE loss exhibits detection rate degradation shortly after a few training epochs (indicating overfitting), the preference-aware loss effectively mitigates overfitting, enabling consistent performance improvement throughout more training epochs.
@@
IV-D Ablation Study
IV-D1 Backbone Architectures

### v5

@@ Abstract
Recent advances in pre-trained language models (PLMs) have demonstrated their capabilities in capturing universal knowledge, making them promising for radar signal processing applications. Nevertheless, directly fine-tuning PLMs on radar signals is both computationally expensive and prone to overfitting, particularly in low signal-to-clutter ratio (SCR) environments. In this paper, we propose a fine-tuning framework for PLM-based marine radar target detection. First, we design a lightweight adaptation module, enabling computationally efficient fine-tuning while preserving the pre-trained model’s general knowledge. Second, a novel preference-aware loss is developed to selectively optimize different feature patches based on their online-evaluated learning values, guiding the model to concentrate on those generalizable feature patterns during optimization. Finally, a binary classification head is retrained based on autoencoder network to further enhance detection performance. Experiments on real-world radar data show that the proposed RadarPLM framework yields at least a 6.35% improvement in detection performance over the existing networks under low SCR conditions. Especially, in the small-sample training cases, the proposed RadarPLM also achieves a significant advantage over existing networks owing to the incorporation of the PLM.
@@ I Introduction
Recently, the research community has experienced a paradigm shift from lightweight, task-specific neural networks toward large-scale foundation models. In particular, pre-trained large language models (LLMs), such as ChatGPT and DeepSeek, have demonstrated strong generalization capabilities across a wide range of complex tasks, including long-form text generation, mathematical reasoning, and code synthesis [12]. Motivated by these advances, an increasing body of research has extended these large models beyond the textual domain to non-textual data modalities, giving rise to multimodal large language models. Meanwhile, several recent studies [13, 14] have explored the adaptation of pre-trained large language models for time-series data analysis and wireless-communication signal processing. These works consistently report notable performance improvements over conventional lightweight task-specific models, particularly in scenarios characterized by limited labeled data. Inspired by these developments, this work investigates the feasibility and effectiveness of leveraging PLMs as a powerful backbone network for marine radar signal processing, aiming to enhance model robustness and generalization under challenging data conditions.
However, adapting PLMs to marine radar signal processing remains non-trivial, primarily due to two critical challenges: the need for parameter-efficient fine-tuning and the risk of model overfitting under complex real-world environments. First, the large parameter scale of PLMs necessitates lightweight and efficient fine-tuning strategies. Without such lightweight adaptation, the extensive computational and memory overhead of PLMs fine-tuning would limit their applicability in real-world radar scenarios. Lightweight adaptation strategy requires reducing the number of trainable parameters by several orders of magnitude while preserving prior knowledge embedded in PLMs. Second, directly adapting PLMs to process radar signal features often leads to severe overfitting. This issue arises from the intrinsic characteristics of radar signal features. On the one hand, radar echoes typically suffer from low signal-to-noise ratios due to complex environmental noise and sensor limitations. Besides, in small marine target detection scenarios, weak target returns are frequently submerged by ocean clutter, producing mixed-status segments that bring noisy or anomalous patterns into the radar signal features. On the other hand, different feature patterns exhibit highly imbalanced convergence behaviors during fine-tuning: simple and regular patterns tend to converge rapidly, while more informative but complex patterns remain under-learned. Consequently, previous uniform optimization strategy may lead to severe overfitting, not only to noisy or anomalous patterns but also to overly simplistic ones. This highlights the necessity of a selective fine-tuning strategy that dynamically adjusts the training emphasis based on the learning value of each feature pattern.
@@
To overcome the above issues, we develop a novel framework named RadarPLM (based on GPT2, an advanced open-source PLM). The proposed framework effectively bridges the gap between PLMs and radar target detection tasks with a lightweight adaptation module and a selective training strategy. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Subsequently, a radar target detector is built upon GPT2 and is fine-tuned through a lightweight adaptation module to efficiently align with the specific radar detection task. To mitigate overfitting, we design a preference-aware loss function, inspired by [15]. By introducing a selective training strategy to dynamically adjust the learning weights of different feature patches according to their online-evaluated learning values. Finally, the binary classification head is trained to further improve detection performance.
@@
    We propose a novel framework RadarPLM based on GPT2 model, an advanced open-source PLM, for marine radar signal processing applications. To the best of our knowledge, this is the first work to demonstrate that the recent developed PLMs can serve as a strong starting point for model training and optimization on radar signal features.
@@
    For fine-tuning, we develop a lightweight adaptation module and design a novel selective training strategy, to reduce fine-tuning computational overhead and significantly mitigate the risk of model overfitting.
@@
    Extensive experiments validate that RadarPLM achieves state-of-the-art detection performance on popular marine radar datasets, especially in low SCR and small-sample training cases.
Our preliminary study [16] initially demonstrated the promising prospects of PLM for radar signal prcocessing applications. In this paper, we further expand that work by introducing a lightweight adaptation module and selective training strategy for more robust and generalizable optimization, achieving better performance.
@@ II-A Marine Radar Target Detection Based on Deep Learning
II-B Signal Processing Based on Pre-trained Language Models
Recently, PLMs such as GPT2 and LLaMA2 have demonstrated exceptional capabilities in sequential signal processing and wireless communication tasks. For example, Zhou et al. [13] propose a unified time series analysis framework by fine-tuning frozen GPT2 [21] and achieving SOTA performance on various datasets. Liu et al. [22] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [23] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [24] propose a multi-task PLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM [25] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks.
However, none of these models are specifically devised for radar signal processing applications. In fact, recent studies [26] have raised significant concerns regarding the universal applicability and direct transferability of PLMs to various signal processing tasks. PLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. Especially, in marine radar target detection under low SCR conditions, directly fine-tuning a PLM often leads to severe overfitting issues, limiting its parameter transfer capability. To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
@@ III Proposed Method
(a) Instantaneous Phase (IP)
(b) Doppler Spectrum Entropy (DSE)
(c) Short-Time Fourier Transform Magnitude Spectrum (SMS)
(d) Amplitude (Amp)
(e) Doppler Phase (DP): Target
(f) Doppler Phase (DP): Clutter
Figure 3: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX dataset.
@@ III-D1 Methodology
  • Low Importance (Noisy/Anomalous patterns): If $\alpha\mathcal{L}_{\theta_{{r}}}$ is large (indicating feature token with inherent noisy/anomalous patterns), it effectively dampens $s_{b,k}$, even if $\mathcal{L}_{\theta_{{t}}}$ is large, preventing model overfitting.
@@ III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
However, directly optimizing (16) is computationally prohibitive, as it requires retraining the model for each candidate patch. To enable a practical surrogate, we reformulate this objective using probabilistic decomposition.
B. Derivation of the Surrogate Objective: Applying Bayes’ theorem, the term in (16) can be decomposed as:
@@
Substituting (18) into (17) and taking the logarithm yields:
@@
In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the holdout samples in (16) can be approximated by the following objective:
@@
D. Conclusion: The calculated gap between the target loss $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})$ and the approximated irreducible loss $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})$ (derived from the lightweight reference model) serves as a practical measure of each feature patch’s learning value, which forms the token weights in (14). Accordingly, the loss gap is employed to adaptively reweight the training objective, guiding the model to focus on learning informative and unlearned feature tokens, while significantly reducing the impact of noisy and trivial patterns.
@@ III-E Autoencoder-Based Binary Classification Head Retraining
As shown in Fig. 1 and Section III-A, the PLM-based network produces token-level outputs of shape $[B,K,2]$, which is inconvenient for final decision. To obtain a binary output for each sample, we therefore retrain an autoencoder-based binary classification head on top of high-dimensional feature representation $F^{\mathrm{LN}}$. It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative representation; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector. The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
@@
    Extract IP, DSE, SMS, Amp, and DP features from the observation vector, respectively, see (3), (4), (5),(7), and (9).
@@
    Compute the loss for each feature token of the training samples based on the predicted output probabilities of $\theta_{r}$, see (12).
@@
    Compute the loss for each feature token of the training samples based on the predicted output probabilities of current model, see (13).
@@
    Fine-tune PLM for target detection via lightweight adaptation module and preference-aware loss function, see (14) and (15).
@@
    Train autoencoder-based binary classification head based on the weighted sum of reconstruction loss and CE loss, see (21), (22), and (23).
@@
    Adjust detection threshold $\eta$ for controllable FAR, see (24).
@@ IV Experiments
Label Name File Name Target Distance Bin SCR
1 17 19931107_135603_starea 9 Low
2 25 19931108_213827_starea 7 Low
3 26 19931108_220902_starea 7 Low
4 54 19931111_163625_starea 8 High
5 280 19931118_023604_stareC0000 8 High
6 283 19931118_035737_stareC0000 10 High
7 311 19931118_162658_stareC0000 7 High
8 320 19931118_174259_stareC0000 7 High
@@ IV-A1 Dataset
To ensure sufficient training data, we employ overlapped segmentation following the partition rule in (1), with parameters set to $M=32$ for target cells and $M=128$ for clutter cells. This process generates 4,079 target samples and more than 9,000 clutter samples per dataset. The samples are divided into three groups: (1) a training set using the first 20% observation time for both the target and clutter cells, (2) a validation set covering 20% to 35% of the observation time, and (3) a test set containing the remaining samples.
@@ IV-A4 Implementation Details
The experimentally relevant hyperparameters and details are shown in Table II. For patching, the size of feature patches is set to 48. For the PLM, the smallest version of GPT2 with $F=768$ feature dimension and the first $L=4$ layers of the GPT2 encoder are deployed. The value of $\alpha$ in (14) is set to 0.9. All signal feature extraction operations are performed with MATLAB R2016a software. All network training experiments are conducted on a system equipped with an E5-2695v3 CPU, an NVIDIA 3090Ti GPU, and 64 GB of RAM.
IV-B Average Performance Comparison
@@
IV-C Performance Comparison on Low SCR Data
Owing to the inherent complexity of the marine environment, low-SCR radar echoes are frequently encountered in practical applications. In such scenarios, clutter-induced noise gives rise to noisy and anomalous feature patterns, leading to pronounced performance degradation in existing detection algorithms. To assess model robustness, we evaluate each approach on more challenging low-SCR datasets (IPIX #17, IPIX #25, and IPIX #26), where the SCR typically falls below 5 dB.
As shown in the right section of Table III, while most baselines suffer from severe performance degradation due to strong clutter-induced noise, RadarPLM maintains a remarkable lead. It achieves an average detection rate of 57.58% on low SCR datasets, surpassing the second-best method (ResNet18) by 6.35% and the complex ADN18 method by a significant margin of 10.27%. This result highlights the critical advantage of our method in learning discriminative representation from noisy and anomalous radar signal features.
IV-D Small-sample Training Condition
@@
Figure 7: Performance comparison under small and abundant sample training conditions. The result shows that the RadarPLM approach experiences substantially smaller performance degradation under scarce training sample condition while maintaining the best overall detection performance.
@@
Figure 8: Comparison of detection rates achieved by the proposed preference-aware loss and the conventional cross-entropy (CE) loss under HH polarization across different datasets ((a) IPIX #17, (b) IPIX #25, and (c) IPIX #26) over training epochs. The results demonstrate that while the CE loss exhibits detection rate degradation shortly after a few training epochs (indicating overfitting), the preference-aware loss effectively mitigates overfitting, enabling consistent performance improvement throughout more training epochs.
@@
IV-E Ablation Study
IV-E1 Backbone Architectures